(5.2.17) Assume that when adults with smartphones are randomly​ selected, 45​% use them in meetings or classes. If 5 adult smartphone users are randomly​ selected, find the probability that exactly 2 of them use their smartphones in meetings or classes
The probability is?
1
Expert's answer
2020-11-29T19:11:37-0500

variations to choose 2 users from 5 is "{5 \choose 2}" . chance that 2 users from 5 use their smartphones in meetings or classes is "0.45^2*0.55^3" . Means probability is

"{5 \choose 2}*0.45^2*0.55^3=\frac {5!} {2!3!}*0.45^2*0.55^3=0.3369" or 33.69%